# PatentMind: A Multi-Aspect Reasoning Graph for Patent Similarity Evaluation

### v1

@@ Abstract
Patent similarity evaluation plays a critical role in intellectual property analysis. However, existing methods often overlook the intricate structure of patent documents, which integrate technical specifications, legal boundaries, and application contexts. We introduce PatentMind, a novel framework for patent similarity assessment based on a Multi-Aspect Reasoning Graph (MARG). PatentMind decomposes patents into three core dimensions: technical feature, application domain, and claim scope, to compute dimension-specific similarity scores. These scores are dynamically weighted through a four-stage reasoning process which integrates contextual signals to emulate expert-level judgment. To support evaluation, we construct PatentSimBench, a human-annotated benchmark comprising 500 patent pairs. Experimental results demonstrate that PatentMind achieves a strong correlation ($r=0.938$) with expert annotations, significantly outperforming embedding-based models and advanced prompt engineering methods.These results highlight the effectiveness of modular reasoning frameworks in overcoming key limitations of embedding-based methods for analyzing patent similarity.
@@ 1 Introduction
Patent documents pose significant challenges for NLP-based similarity evaluation due to specialized domain knowledge, intricate legal language, and complex structural formats (Aslanyan and Wetherbee, 2022; Kanade and Fox, 2007; Casola and Lavelli, 2021; Lupu and Hanbury, 2011). Accurate evaluation of patent similarity is crucial for identifying prior art, assessing infringement risks, and determining the novelty of innovations (Feng, 2020; Helmers et al., 2019; Arts et al., 2018; Wang et al., 2023). Given that inaccuracies can lead to overlooked prior art, invalid patent grants, or undetected infringements, robust and interpretable evaluation methods are essential. Moreover, the rapidly growing volume of patent applications across diverse technological fields further intensifies the need for reliable similarity evaluation techniques (Jiang and Goetz, 2024). Given the increasing reliance on patent analytics for R&D strategy, legal risk assessment, and competitive intelligence, developing accurate and interpretable similarity models has become critically important in both industry and academia (Hain et al., 2022; Lu et al., 2020; Yoo et al., 2023).
@@
Early keyword-based methods, such as bag-of-words models, often fail to capture nuanced semantics and technical specificity of patent documents, resulting in ineffective representations (D’hondt et al., 2013). To overcome these semantic limitations, recent approaches predominantly rely on embedding-based methods, including static word embeddings and transformer-based contextual embeddings (Lee and Hsiang, 2020; Reimers and Gurevych, 2019). However, embedding-based methods compress patent texts into dense vectors, which can obscure critical distinctions and lack interpretability. In contrast, our approach explicitly models similarity through structured reasoning across technical, legal, and application-specific dimensions, closely reflecting how human experts evaluate patent proximity. Such interpretability and precision are particularly critical in legal or technical decision-making contexts (Miric et al., 2023; Bekamiri et al., 2022).
@@
Recent advances in large language models (LLMs) have demonstrated considerable promise in reasoning-based tasks, offering a compelling alternative to traditional embedding-based approaches. Unlike fixed embedding methods, LLMs support dynamic reasoning and nuanced semantic comprehension, facilitating more sophisticated similarity assessments. However, directly applying general-purpose LLMs to patent similarity tasks remains challenging due to specialized terminology and complex features inherent in patent texts (Ikoma and Mitamura, 2025; Ascione and Sterzi, 2024). Moreover, these models often lack domain-specific legal and technical expertise necessary for accurate analysis. Although recent prompting methods for LLMs have partially mitigated the limitations of static embeddings by enabling dynamic contextual analysis, they continue to struggle with domain-specific terminology and processing key patent sections such as titles, abstracts, and claims.
@@
In this work, we propose PatentMind, a novel patent similarity evaluation framework structured as a Multi-Aspect Reasoning Graph (MARG). By leveraging LLM-based multi-step reasoning, PatentMind performs dimension-aware similarity assessments, closely emulating the analytical processes of patent attorneys. Specifically, MARG decomposes patent documents into three essential dimensions: technical features, application domains, and claim scopes. Each dimension is independently analyzed to generate similarity scores, which are then integrated using a context-aware dynamic weighting module consisting of four reasoning stages: Domain Relationship Analysis, Information Distribution, Dimension Relevance, and Cross-Validation. This multidimensional approach mirrors real-world patent examination practices, which evaluate technical details, application contexts, and legal claims in accordance with established standards for prior art analysis (USPTO, 2023). Consequently, PatentMind effectively captures subtle distinctions across the diverse attributes of patents. Our empirical evaluations demonstrate that PatentMind, implemented with GPT-4o-mini as the underlying LLM, achieves a Pearson correlation of 0.938 with expert evaluations, significantly surpassing embedding and LLM-based prompting baselines. To facilitate rigorous evaluation, we introduce PatentSimBench as the first expert-annotated patent similarity benchmark that reflects real-world criteria employed by patent examiners and attorneys. By explicitly modeling technical, legal, and contextual aspects through interpretable reasoning, PatentSimBench and PatentMind collectively address key limitations of prior approaches.
The key contributions in our research PatentMind are summarized as follows:
@@
  • PatentSimBench Dataset: We release the first publicly available expert-annotated benchmark for patent similarity evaluation, comprising 500 high-quality patent pairs. In contrast to existing datasets lacking expert annotations, PatentSimBench provides rigorously annotated pairs with explicit expert rationales, ensuring robust evaluation and reproducibility.
@@ 2.2 Prompt Engineering
2.3 Core Concepts in Patent Documents
To facilitate clarity, we briefly introduce core patent terminology used throughout this paper; detailed definitions are provided in Appendix A. A typical patent document comprises a title, abstract, detailed description and a set of claims that delineate the boundaries of legal protection. The International Patent Classification (IPC) system categorizes patents into technological domains using alphanumeric codes. Key legal concepts include prior art, which refers to existing knowledge that may affect the novelty of an invention, and infringement, which denotes unauthorized use of protected claims. In this work, we define patent similarity across three dimensions: (1) technical attributes: what the invention is; (2) application contexts: where the invention applies; and (3) legal boundaries: how the invention is claimed. These dimensions serve as the foundation for our multi-dimension similarity framework (USPTO, 2020).
@@ 3.1 Multi-Dimensional Feature Extraction
PatentMind captures the complex structure and domain-specific nuances of patent documents through LLM-based feature extraction. We decompose each patent into three dimensions: technical feature ($T$), application domain ($D$), and claim scope ($C$). These dimensions reflect how patent specialists evaluate patent similarity, focusing on what the invention is, where it applies, and how broadly it is claimed (USPTO, 2023).
@@ 3.3 Context-Aware Dynamic Weight Reasoning
Figure 3: The computational graph of Context-Aware Dynamic Weight Reasoning framework.
PatentMind adapts to contextual nuances by computing weights for each similarity dimension through multi-step LLM reasoning. This process comprises four key stages.
@@ 4 Dataset Construction
To facilitate rigorous evaluation of patent similarity methods, we introduce PatentSimBench, the first expert-annotated benchmark dataset tailored for patent similarity tasks. PatentSimBench provides rationale-supported similarity scores aligned with real-world legal and technical judgments, making it particularly suitable for evaluating multi-step reasoning models. The dataset consists of 500 patent pairs sampled from the USPTO database (USPTO, 2023), with a balanced representation across all IPC sections. The International Patent Classification (IPC) categorizes patents into eight top-level domains, such as human necessities, chemistry, physics, and digital technologies (WIPO, 2023), ensuring comprehensive technical coverage across diverse fields. Similarity annotations are based on a 5-point Likert scale, with scores ranging from 1 (minimal similarity) to 5 (maximal similarity). Our dataset is publicly available on GitHub (It will be available upon acceptance.).
@@ 4.2 Annotation Process
4.3 Comparison with Existing Datasets
PatentSimBench is the first expert-annotated patent similarity dataset. Although the dataset comprises only 500 pairs, the size is a deliberate design choice given the legal and technical complexity of patent documents, which makes large-scale expert annotation impractical. Instead, PatentSimBench prioritizes high annotation quality and technical diversity, capturing nuances of real-world similarity evaluations that automated methods struggle to reflect. Annotated by four experts through a rigorous consensus process, it offers a reliable benchmark with sufficient scale and depth to robustly evaluate AI-driven patent similarity models. Its strength lies not in size, but in qualitative excellence–setting a new standard for research in patent analysis.
@@ 5 Experimental Results
5.1 Comparison with Embedding-Based Methods
@@
PatentMind .938 .923 .013 .092
Table 2: The performance comparison of PatentMind with varying embedding-based baseline methods.
Table 2 compares PatentMind with conventional embedding-based models. Baseline approaches show only moderate correlation with expert judgments with high error rates. In contrast, PatentMind achieves significantly better performance with a Pearson correlation of 0.938 and the lowest error rates, demonstrating superior alignment with human evaluations.
@@ 5.2 Comparison with Prompting-Based Methods
[…]
6 Ablation Studies
@@ 6.2 Comparison of Regression Weighting
Fold Type Pearson Spearman RSME MAE
@@
[…]
@@ 6.3 Comparison of Aspect Impact
To assess the contribution of each component in our framework, we conducted an ablation study across seven configurations, including the removal of dynamic weighting, exclusion of individual dimensions, and single-dimension-only variants. As shown in Table 6, replacing the context-aware dynamic weighting mechanism with equal weighting results in a noticeable performance drop (3.35%), demonstrating the importance of adapting weights to contextual relevance. Removing any one of the three core dimensions(Claim Scope, Technical, or Application)also led to performance degradation (average drop: 2.82%), indicating their complementary roles in similarity evaluation. Notably, the removal of Technical (3.10%) or Application (3.20%) dimensions caused greater performance loss than removing Claim Scope (2.15%). Single-dimension variants showed even larger drops in performance (average 4.03%), with the Application-only model suffering the most (7.40%). Although the Technical-only model performed relatively better (2.05% drop), it still fell short of the full model’s performance. These findings validate the effectiveness of our design choices: the integration of technical, legal, and domain-level reasoning, combined with context-aware dynamic weighting, leads to substantially improved accuracy and adaptability in patent similarity evaluation.
@@ 7 Error Analysis
In this section, we examine the performance boundaries of PatentMind with GPT-4o-mini as the backbone LLM. We analyze prediction accuracy and residual patterns, identify domain-specific error trends across IPC sections, and investigate high-error cases involving functional ambiguity and terminology mismatch. This analysis reveals the model’s strengths and limitations and informs concrete directions for future improvement.
@@ 7.1 Prediction and Residual Analysis
We assess PatentMind’s reliability by comparing predicted similarity scores to expert annotations and analyzing residuals. Predictions show strong alignment, with most errors within the 0–0.1 range (MAE = 0.0918). Figure 4 shows predictions generally follow the ideal ($y=x$) line, though high-similarity pairs (ground truth $>0.8$) tend to be slightly underestimated, and low-similarity pairs ($<0.2$) slightly overestimated. The average residual ($-0.0197$) indicates minimal bias, with domain-specific variations discussed further below.
7.2 Error Analysis by IPC Code
Our IPC-level analysis shows that PatentMind’s prediction errors are more pronounced in certain technical domains, such as metallurgy (C22B), genetic engineering (C12N), and mechanical engineering (F16D). These discrepancies stem from (1) domain-specific terminology underrepresented in LLM pretraining, (2) structural variation in patent writing across fields, and (3) similarity criteria. For instance, metallurgy patents (e.g., C22B300) often suffer from underestimated similarity due to synonym recognition issues (e.g., “leaching” vs. “extraction”), while audio processing patents (e.g., G06F316) tend to be overestimated due to generic language. Biotechnology patents (C12N) present additional challenges due to linguistic complexity and causal reasoning structures.
These insights highlight the need to incorporate domain-specific knowledge into patent similarity assessment. While PatentMind’s dynamic weighting indirectly reflects IPC-related contexts, it lacks explicit modeling of specialized expertise. However, its modular design enables integration of domain-adaptive modules without major redesign, a key direction for future improvement.
@@ 8 Conclusion
PatentMind is a novel framework for patent similarity evaluation that structurally analyzes patents across three core dimensions: technical features, application domains, and claim scope. By dynamically aggregating aspect-wise similarity scores through LLM-based contextual reasoning, the framework achieves high alignment with expert judgments and outperforms traditional embedding-based methods in both accuracy and interpretability. Notably, this work addresses the unique challenges posed by patents as hybrid legal–technical documents, proposing a comprehensive evaluation framework that integrates the linguistic precision of NLP, the normative criteria of legal reasoning, and the structural complexity of technical content. These results demonstrate the practical value and extensibility of semantically grounded, structured approaches in legal NLP.
@@
Dataset Scale and Diversity.
The PatentSimBench dataset consists of 500 patent pairs, constructed through an expert annotation process by 4 domain experts, achieving high consistency as measured by Fleiss’ Kappa (0.588) and Cronbach’s Alpha (0.967). This dataset is publicly available, providing a valuable resource for subsequent research. While it effectively balances IPC technical fields for initial validation, its scope remains limited relative to expansive patent repositories. In future work, we plan to expand the dataset’s scale and incorporate greater domain diversity to enhance generalizability.
@@
The multi-stage reasoning structure and dynamic weight calculation in PatentMind provide high precision and interpretability but introduce relatively higher computational overhead. This could potentially constrain deployment in large-scale real-time applications; however, we consider this a necessary trade-off in tasks like patent analysis where expertise and accuracy are paramount. Our experimental results demonstrate consistency with expert evaluations, validating this approach. Subsequent research will explore model optimization techniques and efficient prompt designs to improve computational efficiency and enhance scalability.
These limitations do not diminish the major contributions of PatentMind. Rather, they reinforce the sustainability and advancement potential of our research. By explicitly acknowledging these constraints and proposing specific directions for improvement, our work establishes long-term impact potential in the field of patent similarity evaluation.
Domain-Specific Limitations.
While PatentMind demonstrates strong overall performance, our error analysis reveals varying effectiveness across technical domains. The current framework employs a general approach that may not optimally capture domain-specific nuances in highly specialized fields. As shown in Section  7.2, domains with complex terminology (e.g., metallurgy, biotechnology) or unique structural conventions exhibit higher error rates. Although our dynamic weighting mechanism partially addresses this by considering IPC code relationships, it lacks explicit domain-specific knowledge integration. Furthermore, our current evaluation, while covering diverse IPC sections, is limited in its ability to provide statistically significant insights at more granular domain levels due to dataset size constraints. Future work should explore more targeted domain adaptation techniques, including specialized prompts for high-error domains, domain-specific terminology processing, and potential fine-tuning approaches for technical fields where general LLMs show limitations in capturing specialized knowledge. Importantly, the modular design of PatentMind allows for seamless integration of domain-adaptive components without reengineering the overall architecture.
@@
This research adheres to ethical principles in data utilization. The PatentSimBench dataset comprises exclusively publicly available patent documents, complying with intellectual property regulations and containing no personally identifiable information. Acknowledging potential biases in language models, we implemented expert annotations and structured reasoning methodologies to enhance the interpretability and fairness of the results. To facilitate transparency and reproducibility, we have made our research methodology and dataset publicly accessible. We have considered the potential misuse of AI-based patent similarity evaluation tools in legal disputes or competitive intelligence analysis. Consequently, we recommend using this system as a supplementary tool rather than as the sole basis for high-stakes decisions. Our future research will continue investigating bias mitigation strategies and enhancing ethical applications of patent evaluation systems.
Acknowledgements
We extend our sincere gratitude to collaborators whose insightful discussions and feedback substantially enhanced this research. Special thanks to the expert annotators who contributed to the PatentSimBench dataset, ensuring robust and high-quality patent similarity evaluations. We are grateful to our affiliated institutions and funding agencies for providing essential research resources. We also thank OpenAI, Meta, Alibaba and Anthropic for access to their large language models, which enabled comprehensive cross-LLM evaluation of our framework.

### v2

@@ Abstract
Patent similarity evaluation plays a critical role in intellectual property analysis. However, existing methods often overlook the intricate structure of patent documents, which integrate technical specifications, legal boundaries, and application contexts. We introduce PatentMind, a novel framework for patent similarity assessment based on a Multi-Aspect Reasoning Graph (MARG). PatentMind decomposes patents into their three dimensions of technical features, application domains, and claim scopes, then dimension-specific similarity scores are calculated over the MARG. These scores are dynamically weighted through a context-aware reasoning process, which integrates contextual signals to emulate expert-level judgment. To support evaluation, we construct a human-annotated benchmark PatentSimBench, comprising 500 patent pairs. Experimental results demonstrate that the PatentMind-generated scores show a strong correlation ($r=0.938$) with expert annotations, significantly outperforming embedding-based models, patent-specific models, and advanced prompt engineering methods. Beyond computational linguistics, our framework provides a structured and semantically grounded foundation for real-world decision-making, particularly for tasks such as infringement risk assessment, underscoring its broader impact on both patent analytics and evaluation.
@@ 1 Introduction
Patent documents pose significant challenges for NLP-based similarity evaluation due to specialized domain knowledge, intricate legal language, and complex structural formats (Aslanyan and Wetherbee, 2022; Indukuri et al., 2007; Casola and Lavelli, 2021; Lupu and Hanbury, 2013). Accurate evaluation of patent similarity is crucial for identifying prior art, assessing infringement risks, and determining the novelty of innovations (Feng, 2020; Helmers et al., 2019; Arts et al., 2018; Wang et al., 2023). Given that inaccuracies can lead to overlooked prior art, invalid patent grants, or undetected infringements, robust and interpretable evaluation methods are essential. Moreover, the rapidly growing volume of patent applications across diverse technological fields further intensifies the need for reliable similarity evaluation techniques (Jiang and Goetz, 2024). Given the increasing reliance on patent analytics for R&D strategy, legal risk assessment, and competitive intelligence, developing accurate and interpretable similarity models has become critically important in both industry and academia (Hain et al., 2022; Yoo et al., 2023).
@@
In the early stages of patent similarity evaluation, keyword-based methods such as bag-of-words were widely used. However, these approaches struggled to capture the semantics and technical specificity of patent documents, resulting in ineffective representations (D’hondt et al., 2013; Ascione and Sterzi, 2024). To overcome these limitations, embedding-based methods such as static word embeddings and transformer-based contextual embeddings were subsequently introduced (Lee and Hsiang, 2020; Reimers and Gurevych, 2019). Nevertheless, these methods compress patent texts into dense vectors, which can obscure critical distinctions and reduce interpretability.
@@
Recent advances in large language models (LLMs) have demonstrated considerable promise in reasoning-based tasks, offering a compelling alternative to traditional embedding-based approaches. Unlike fixed embedding methods, LLMs support dynamic reasoning and semantic comprehension, facilitating more sophisticated similarity assessments. However, directly applying general-purpose LLMs to patent similarity tasks remains challenging due to specialized terminology and complex features inherent in patent texts (Ikoma and Mitamura, 2025; Ascione and Sterzi, 2024). Moreover, prompt-based approaches to LLMs have partially addressed the limitations of embeddings, but still struggle due to the specialized terminology and structural complexity of patent documents.
To address these challenges, we propose PatentMind, a novel patent similarity evaluation framework structured as a Multi-Aspect Reasoning Graph (MARG). This framework emulates the analytical processes of patent experts through multi-step reasoning based on LLMs. PatentMind decomposes patent documents into three core dimensions: technical features, application domains, and claim scopes. Each dimension is independently evaluated for similarity, and the results are integrated through a context-aware reasoning process consisting of four stages: Domain Relationship Analysis, Information Distribution, Dimension Relevance, and Cross-Validation. This multi-dimensional approach is designed to closely reflect real-world patent examination practices. (USPTO, 2024).
@@
Consequently, PatentMind effectively captures subtle distinctions across the diverse attributes of patents. Our empirical evaluations demonstrate that PatentMind, implemented with GPT-4o-mini as the underlying LLM, achieves a Pearson correlation of 0.938 with expert evaluations, significantly surpassing embedding and LLM-based prompting baselines. To facilitate rigorous evaluation, we introduce PatentSimBench as the first expert-annotated patent similarity benchmark that reflects real-world criteria employed by patent examiners and attorneys. By explicitly modeling technical, legal, and contextual aspects through interpretable reasoning, PatentMind collectively address key limitations of prior approaches.
The key contributions of PatentMind are summarized as follows:
@@
  • Performance Superiority: We demonstrate that PatentMind consistently surpasses embedding-based, prompting-based, and even fine-tuned models that require substantial time and computational resources, with this advantage arising from its patent-specialized, multi-dimensional reasoning framework rather than mere reliance on LLM calls.
  • PatentSimBench Dataset: We release the first publicly available expert-annotated benchmark for patent similarity evaluation, comprising 500 high-quality patent pairs.
@@ 3.1 Multi-Dimensional Feature Extraction
PatentMind captures the complex structure and domain-specific nuances of patent documents through LLM-based feature extraction. We decompose each patent into three dimensions: technical feature ($T$), application domain ($D$), and claim scope ($C$). These dimensions reflect how patent specialists evaluate patent similarity, focusing on what the invention is, where it applies, and how broadly it is claimed (USPTO, 2024).
@@
Figure 3: The computational graph of the Context-Aware Dynamic Weight Reasoning framework, with modules annotated by their notations: Domain Relationship Analysis ($f_{rel},R_{domain}$), Information Distribution Analysis ($f_{dist},D_{info}$), Dimension Relevance Assessment ($f_{assess},A_{rel}$), and Cross-validation Reasoning ($f_{valid},V_{cross}$).
@@ 3.3 Context-Aware Dynamic Weight Reasoning
PatentMind adapts to contextual nuances by computing weights through multi-step LLM reasoning. Instead of parameter-intensive fine-tuning, it functions as an open calculation process using reasoning prompts, highlighting both efficiency and adaptability. This process comprises four key stages.
@@ 4 Dataset Construction
To facilitate rigorous evaluation of patent similarity methods, we introduce PatentSimBench, the first expert-annotated benchmark dataset tailored for patent similarity tasks. PatentSimBench provides rationale-supported similarity scores aligned with real-world legal and technical judgments, making it particularly suitable for evaluating multi-step reasoning models. The dataset consists of 500 patent pairs sampled from the USPTO database (USPTO, 2025), with a balanced representation across all IPC sections. The International Patent Classification (IPC) categorizes patents into eight top-level domains, such as human necessities, chemistry, physics, and digital technologies (WIPO, 1975), ensuring comprehensive technical coverage across diverse fields. Similarity annotations are based on a 5-point Likert scale, with scores ranging from 1 to 5. Our dataset is publicly available on GitHub(It will be available upon accept).
@@ 5 Experimental Results
5.1 Comparison with Embedding-Based and Patent specific Methods
@@
Patent-GPT-J .892 .883 .124 .178
PatentMind .938 .923 .113 .092
Table 2: The performance comparison of PatentMind with varying embedding-based and patent specific baseline methods.
Table 2 compares PatentMind with conventional embedding-based and patent-specific models. Baseline approaches exhibit only moderate correlation with expert judgments and show high error rates. In contrast, PatentMind achieves significantly better performance with a Pearson correlation of 0.938 and the lowest error rates, demonstrating superior alignment with human evaluations.
@@ 5.2 Comparison with Prompting-Based Methods
We compare PatentMind with representative prompting strategies in Table 3. While advanced methods such as Few-shot, Chain-of-Thought (CoT), and Self-Consistency yield reasonable performance, they struggle to capture the multi-dimensional nature of patent documents. For instance, CoT typically follows a single reasoning path, limiting its ability to jointly model technical, legal, and domain-specific aspects. Self-Consistency introduces variation in reasoning but fails to incorporate context-sensitive judgment aligned with expert reasoning. In contrast, PatentMind integrates hierarchical reasoning with dynamic weighting, enabling more faithful modeling of patent semantics. As a result, it achieves the highest agreement with expert judgments (Pearson $r=.938$; Spearman $\rho=.923$). These findings highlight the critical role of explicitly modeling multi-aspect patent contexts in enhancing both accuracy and interpretability.
@@ 5.4 Comparison of Regression Weighting
Fold Type Pearson Spearman RMSE MAE
@@
To validate the effectiveness of PatentMind’s dynamic weighting strategy, we compared it against several regression-based baselines using the PatentSimBench dataset. Model performance was evaluated based on Pearson and Spearman correlations between predicted similarity scores and expert annotations. As shown in Table 5, all regression models, including Linear Regression, Lasso, SVM, Tree Boosting, Bayesian Ridge, and a 2-layer MLP Regressor, underperformed relative to PatentMind. The best-performing baseline, MLP Regression (Pearson $r=.921$; Spearman $\rho=.896$), still fell short of PatentMind’s performance (Pearson $r=.938$; Spearman $\rho=.923$). These results reveal that regression-based fixed-weighting methods struggle to capture the nuanced, context-sensitive variations essential for accurate patent similarity evaluation. Unlike these baselines, which require supervised training to learn fixed parameters, PatentMind functions as an open reasoning process driven by LLM prompts, avoiding extra training costs while dynamically adapting weights to each patent pair.
6 Ablation Studies
@@ 6.1 Comparison of Aspect Impact
To assess the contribution of each component in our framework, we conducted an ablation study across seven configurations, including the removal of dynamic weighting, exclusion of individual dimensions, and single-dimension-only variants. As shown in Table 6, replacing the context-aware dynamic weighting mechanism with equal weighting results in a noticeable performance drop (3.35%), demonstrating the importance of adapting weights to contextual relevance. Removing any one of the three core dimensions(Claim Scope, Technical, or Application)also led to performance degradation (avg. 2.82%), indicating their complementary roles in similarity evaluation. Notably, the removal of Technical (3.10%) or Application (3.20%) dimensions caused greater performance loss than removing Claim Scope (2.15%). Single-dimension variants showed even larger drops in performance (avg. 4.03%), with the Application-only model suffering the most (7.40%). Although the Technical-only model performed relatively better (2.05% drop), it still fell short of the full model’s performance. These findings validate the effectiveness of our design choices: the integration of technical, legal, and domain-level reasoning, combined with context-aware dynamic weighting, leads to substantially improved accuracy and adaptability in patent similarity evaluation.
@@ 7 Error Analysis
In this section, we examine the performance boundaries of PatentMind. We analyze prediction accuracy and residual patterns, identify domain-specific error trends across IPC sections, and investigate high-error cases involving functional ambiguity and terminology mismatch.
@@ 7.1 Prediction and Residual Analysis
We assess PatentMind’s reliability by comparing predicted similarity scores to expert annotations and analyzing residuals. Predictions show strong alignment, with most errors within the 0–0.1 range (MAE = 0.0918). Figure 4 shows predictions generally follow the ideal ($y=x$) line, though high-similarity pairs (ground truth >0.8>0.8) tend to be slightly underestimated, and low-similarity pairs (<0.2<0.2) slightly overestimated. The average residual (0.0197-0.0197) indicates minimal bias, with domain-specific variations discussed further below.
7.2 Error Analysis across Different IPC Codes
Our IPC-level analysis shows that PatentMind’s prediction errors are more pronounced in certain technical domains, such as metallurgy (C22B), genetic engineering (C12N), and mechanical engineering (F16D). These errors stem from: (1) underrepresentation of domain-specific terms in LLM pretraining (e.g., “leaching” vs. “extraction” in C22B), (2) structural variations in patent drafting across domains, and (3) context-dependent similarity criteria. For instance, metallurgy patents (e.g., C22B300) often suffer from underestimated similarity due to synonym recognition issues (e.g., “leaching” vs. “extraction”), while audio processing patents (e.g., G06F316) tend to be overestimated due to language. Biotechnology patents (C12N) present additional challenges due to linguistic complexity and causal reasoning structures.
These findings underscore the need for domain-adaptive embeddings or knowledge graph integration to enhance synonym recognition and contextual understanding. PatentMind’s modular design facilitates such enhancements without requiring extensive retraining, a key direction for future work.
@@ 8 Conclusion
In this paper, we proposed PatentMind, a novel framework for patent similarity evaluation that integrates linguistic, legal, and technical reasoning.
First, we introduced a multi-dimensional analysis that, for the first time, jointly considers the three core aspects of patents: technical features, application domains, and claim scope. This approach effectively captures the structural complexity of patent documents.
Second, we proposed context-aware weighting, where LLM-based reasoning dynamically adjusts the relative importance of each dimension depending on domain- and case-specific contexts. This is the first method to discriminatively capture key factors under varying patent contexts.
Third, we achieved expert-level agreement and demonstrated clear superiority over embedding-based and prompting-based baselines, showing that our approach goes beyond simple LLM calls or prompting strategies.
Fourth, we showed that PatentMind outperforms specialized patent-tuned models, demonstrating higher accuracy while being more efficient than fine-tuning approaches that require substantial time and computational resources.
Fifth, we established the model-agnostic robustness of PatentMind. It consistently delivers strong performance across diverse LLMs, ensuring generalizability and reproducibility without dependence on any single vendor or architecture.
Finally, we introduced PatentSimBench, the first expert-annotated benchmark for patent similarity, which provides a valuable resource for both academia and industry.
Taken together, these contributions extend beyond computational linguistics, offering a structured and semantically grounded foundation for addressing real-world decision-making tasks such as prior art search, infringement risk assessment, and novelty evaluation.
@@
PatentMind’s multi-stage reasoning structure and dynamic weighting mechanism deliver high precision and interpretability but introduce relatively high computational complexity, which may necessitate additional optimization for large-scale or latency-sensitive applications. This issue is distinct from PatentMind’s core research objective (emphasizing accuracy in patent similarity evaluation) and will be addressed in future work. Specifically, subsequent efforts will explore optimizations such as response caching, selective execution of reasoning steps, and lightweight model substitution (e.g., knowledge distillation or LoRA-based LLMs). These approaches are informed by recent studies demonstrating that LLM distillation techniques, such as the Distilling Step-by-Step method, can effectively transfer multi-stage reasoning capabilities to smaller models, and that LoRA’s low-rank adaptation is well-suited for dynamically optimizing model weights. Furthermore, PatentMind’s modular design enables flexible adaptation to efficiency requirements, and techniques like prompt compression and batch parallel processing (e.g., tensor parallelization or in-flight batching for optimized GPU utilization) can further enhance scalability in deployment environments.
These limitations do not diminish PatentMind’s primary contributions, including achieving a 0.938 Pearson correlation with expert annotations. Rather, they underscore the research’s sustainability and potential for advancement. By explicitly acknowledging these constraints and proposing concrete directions for improvement, this work establishes a long-term impact in the field of patent similarity evaluation.
Domain-Specific Opportunities.
While PatentMind demonstrates strong overall performance, our error analysis highlights opportunities for further enhancement across different technical domains. The current framework adopts a general approach, which provides broad applicability but can be further strengthened to capture domain-specific nuances in highly specialized fields. As shown in Section  7.2, domains with complex terminology (e.g., metallurgy, biotechnology) or unique structural conventions represent promising areas where targeted improvements may yield additional gains.
Our dynamic weighting mechanism already helps address these challenges by considering IPC code relationships, and future work can build on this by integrating explicit domain knowledge. Although the current evaluation spans diverse IPC sections, expanding dataset coverage will enable more fine-grained statistical insights at specific domain levels. This creates an opportunity to design more specialized adaptation strategies.
In particular, tailored prompts for high-error domains, domain-specific terminology processing, and fine-tuning approaches for technical fields can further amplify PatentMind’s effectiveness. Importantly, the modular design of PatentMind ensures that such domain-adaptive components can be integrated seamlessly without reengineering the overall architecture, underscoring its flexibility and extensibility for future advancements.
@@
This research adheres to ethical principles in data utilization. The PatentSimBench dataset comprises exclusively publicly available patent documents, complying with intellectual property regulations and containing no personally identifiable information. Acknowledging potential biases in language models, we implemented expert annotations and structured reasoning methodologies to enhance the interpretability and fairness of the results. To facilitate transparency and reproducibility, we have made our research methodology and dataset publicly accessible. We have considered the potential misuse of AI-based patent similarity evaluation tools in legal disputes or competitive intelligence analysis. Consequently, we recommend using this system as a supplementary tool rather than as the sole basis for high-stakes decisions. Our future research will continue investigating bias mitigation strategies and enhancing ethical applications of patent evaluation systems. For transparency, we note that ChatGPT was used to improve grammar and clarity of expression.